# CoT-Self-Instruct: Building high-quality synthetic prompts for reasoning and non-reasoning tasks

### v1

@@ 4.2 Non-verifiable instruction following
We use the Wildchat-RIP-Filtered-by-8b-Llama dataset [^2] https://huggingface.co/datasets/facebook/Wildchat-RIP-Filtered-by-8b-Llama which includes 4k high-quality prompts filtered from 20k raw wildchat prompts as our seed prompts. Unlike in some reasoning tasks, this data includes a variety of different domains. Combining, for instance, prompts from storytelling and coding as few-shot examples could result in generating unnatural synthetic prompts. To address this, we categorized all seed data into 8 distinct categories. During sampling, we select 2 seed prompts from the same category to serve as few-shot prompts. Our seed data spans 8 categories: Writing & Storytelling, Technical & Programming, Creative & Design, Data & Analysis, Education & Research, Communication & Support, Business & Marketing, and Miscellaneous.
Prompt Generation
We evaluate how CoT-Self-Instruct compares to baselines for generating non-verifiable instruction-following tasks by applying these methods using LLama 3.1-8B-Instruct. The CoT-Self-Instruct template is given in Figure 3, and the baseline Self-Instruct method is given in Figure 4. We also experiment with a prompt that lies between the two methods by generating a short rather than long CoT, given in Figure 5. For RIP filtering, we use the Athene-RM-8B model [^3] https://huggingface.co/Nexusflow/Athene-RM-8B. reward model over 32 responses.
@@
We train via DPO starting from LLama 3.1-8B-Instruct, leveraging the fairseq2 library (Balioglu, 2023). We use a batch size of 646464 and learning rate of $1e{-6}$ with dropout rate of 0.0 and a $\beta$ value of 0.1 throughout the experiments. For each prompt, we generate 64 responses. These responses are then annotated with Athene-RM-8B to select pairs. Compared to human prompts, our synthetic prompts tend to be more complex, resulting in longer average response lengths, which can lead to length explosion. During DPO training, the evaluation judge often favors longer responses, potentially causing response lengths to increase over time (Yuan et al., 2024). To mitigate this issue, we adopted the approach outlined by Wu et al. (2024), which involves combining the reward score with length information to determine the preferred response. This method ensures that shorter responses are selected when scores are similar. We applied a length normalization coefficient of 0.2 for the length-normalized reward. This is applied for all methods, in each case sampling 5k DPO pairs.
@@
We also assess online DPO by following the training settings described by Lanchantin et al. (2025). We use the default sampling parameters (temperature=1.01.01.0, top-p=1.01.01.0) to generate exploration rollouts. We train models using the fairseq2 library (Balioglu, 2023), where model inference is performed with the vllm library (Kwon et al., 2023).
@@
To evaluate the helpfulness and quality of responses, we employ AlpacaEval 2.0 (Li et al., 2023b; Dubois et al., 2024) and Arena-Hard (Li et al., 2024b; a). These are robust instruction-following benchmarks that show a strong correlation with user preferences. Originally, AlpacaEval used GPT-4 Preview (11/06) as the judge, while Arena-Hard utilized GPT-4.1 for its leaderboard. However, since we do not have access to these specific OpenAI API versions, we conduct our tests using two alternative (and newer) judges: GPT-4-turbo and GPT-4o. For generating predictions, we set the decoding temperature to 0.6 and the top-p to 0.9, aligning with the commonly used values of the seed model in our study. Our validation set used for checkpoint selection is based on a held-out set of 470 examples, comprising 253 validation examples from Li et al. (2023a) and 218 Evol-Test set examples from Xu et al. (2023).
@@ 5.1 Reasoning Tasks
In Table 1 where Qwen3-4B-Base models are GRPO trained on Qwen3-4B generated prompts and responses, CoT-Self-Instruct achieves an average accuracy of 53.0%, outperforming Self-Instruct which yields 49.5% (both without data filtering). As shown in Appendix Table 8 similar trends are observed when training on Qwen3-4B-base, rather than Qwen3-4B, generated targets.
@@
Applying filtering methods to both CoT-Self-Instruct and Self-Instruct improves both methods, despite the overall amount of training data decreasing, see Table 1. That is, it is better to have high quality synthetic data than more data that is lower quality. However, we find that CoT-Self-Instruct still maintains its advantage over Self-Instruct, whichever filtering method is used. E.g. with Self-Consistency Filtering Self-Instruct improves from 49.5% $\rightarrow$ 53.6%, while CoT-Self-Instruct improves from 53.0% $\rightarrow$ 55.1%. Similar findings are observed with RIP filtering as well. We find empirically that the optimal filtering criteria is to filter out prompts with lower than 50% RIP-Scores.
Table 1: CoT-Self-Instruct results on reasoning tasks, comparing to baselines, fine-tuning Qwen3-4B-Base with GRPO. For Self-Instruct and CoT-Self-Instruct the synthetic data (including targets) is constructed with Qwen3-4B. We report pass@1 averaged over 16 seeds. CoT-Self-Instruct generates synthetic data that outperforms existing prompt training sets and the Self-Instruct method, particularly when applying our data filtering methods.
@@
s1k Prompts + (R1) Gold Label 893 68.6 18.5 51.3 40.1 44.6
OpenMathReasoning Prompts + Gold Label 10,000 79.0 13.3 62.5 35.4 47.5
Self-Instruct 5000 81.1 16.3 58.1 42.5 49.5
@@
+ Answer-Consistency Filter (more data) 10,000 86.7 26.7 73.8 47.4 58.7
Table 2: CoT-Self-Instruct results on general instruction following tasks, comparing to baselines, fine-tuning LLama 3.1-8B-Instruct with offline and online DPO. CoT-Self-Instruct generates synthetic data that outperforms human-written prompt training sets and the Self-Instruct method, particularly when applying our data filtering methods. Both AlpacaEval 2 and ArenaHard are evaluated with two kinds of judge: GPT-4 Turbo and GPT-4o, with similar conclusions.
Training AlpacaEval LC Winrate ArenaHard Score
Method GPT-4 Turbo GPT-4o GPT-4 Turbo GPT-4o Avg. $\uparrow$
@@
Human prompts (WildChat) DPO 49.1 43.0 52.7 42.6 46.8
@@
Human prompts (Wildchat) Online DPO 80.1 62.7 64.4 45.5 63.1
@@
High quality synthetic prompts generated by CoT-Self-Instruct significantly outperform seed instructions and other publicly available reasoning prompts
CoT-Self-Instructions outperform s1k, see Table 1, where models trained on 2926 filtered CoT-Self-Instructions achieve 57.2%. This is much higher than the 44.6% achieved with s1k prompts using R1 labels (and s1k results are even lower, 43.8%, with Qwen3-4B labels, see Appendix Table 5). Filtering CoT-Self-Instruction to the same training size as s1k yields 54.2%, still significantly higher, see Appendix Table 3. These results also outperform using 10K OpenMath-Reasoning instructions with gold labels, which gives 47.5%. Increasing the CoT-Self-Instruction with Answer-Consistency filtering data to 10k improves results further with an average of 58.7%. Overall, CoT-Self-Instruct with Answer-Consistency filter gives the best performance of all existing datasets or synthetic data construction methods tested.
5.2 Non-verifiable instruction following tasks
@@
For non-reasoning tasks, allowing the model to create a plan beforehand with CoT-Self-Instruct also significantly enhances the quality of synthetic data, see Table 2. Averaged over AlpacaEval 2 and ArenaHard, CoT-Self-Instruct achieves an average of 53.9 vs. Self-Instruct’s 47.4, both without filtering and training with DPO. We also observe that asking for longer CoT reasoning chains provides more gains that shorter CoTs, see Appendix Table 11, further emphasizing the need for reasoning when producing synthetic data.
@@
Applying the RIP filter to each method we find it to be effective across all types of synthetic generation methods tested. This boosts the CoT-Self-Instruct results from 53.9 $\rightarrow$ 54.7. RIP also improves Self-Instruct as well, from 47.4 $\rightarrow$ 49.1, but still underperforming CoT-Self-Instruct. We can also apply RIP filtering to human prompts from WildChat in a similar manner. In this case we actually see a larger boost, from 46.8 $\rightarrow$ 50.7. We attribute this to human data being relatively noisy compared to synthetic data, which can make filtering more important.
High quality synthetic prompts generated by CoT-Instruct significantly outperform human prompts
Our best performing DPO-trained model is achieved by using CoT-Self-Instruct with RIP data filtering, yielding 54.7. This outperforms LLama 3.1-8B-Instruct (27.1) or training on human prompts from WildChat with (46.8) or without RIP data filtering (50.7). We also performed experiments with online DPO, which improved results further. In that setting human prompts from WildChat obtain 63.1 while CoT-Self-Instruct+RIP obtains 67.1. Overall, we find CoT-Self-Instruct with RIP filtering to yield the best performance over all existing datasets or synthetic data construction methods tested.
@@ 6 Conclusion
In this paper, we propose CoT-Self-Instruct, a synthetic data creation and curation pipeline that instructs LLMs to plan and reason to come up with new synthetic prompts given seed instructions, and then filters them for quality, either using Answer-Consistency for verifiable tasks or RIP filtering when they are not verifiable. We show that applying our method improves models’ abilities in both the reasoning and non-reasoning domains by creating high quality synthetic instructions for RL training, surpassing existing seed human-annotated instructions and public training sets on challenging benchmarks.

### v2

@@ 4.2 General Instruction Following
As our seed examples, we use the Wildchat-RIP-Filtered-by-8b-Llama dataset [^2] https://huggingface.co/datasets/facebook/Wildchat-RIP-Filtered-by-8b-Llama, which includes 4k high-quality instructions filtered from 20k raw Wildchat samples. We prompt the LLama 3.3-70B-Instruct model to classify each seed instruction into one of the following 8 categories: Writing & Storytelling, Technical & Programming, Creative & Design, Data & Analysis, Education & Research, Communication & Support, Business & Marketing, and Miscellaneous.
Instruction generation
We evaluate how CoT-Self-Instruct compares to the baselines for generating non-verifiable instruction-following tasks by training LLama 3.1-8B-Instruct on the generated data. The CoT-Self-Instruct template is given in Figure 3, and the baseline Self-Instruct method is given in Appendix Figure 4. We also experiment with an approach that lies between the two methods by generating a shorter, rather than a longer, CoT, as shown in Appendix Figure 5. We prepare a set of 5,000 few-shot prompts, each consisting of two randomly selected seed examples from the same category. For RIP filtering, we generate 32 synthetic instructions from each few-shot prompt, and keep the one with the highest RIP score, resulting in 5,000 synthetic instructions in total. We use the Athene-RM-8B [^3] https://huggingface.co/Nexusflow/Athene-RM-8B. reward model for the RIP scoring.
@@
We train via DPO starting from Llama 3.1-8B-Instruct, leveraging the fairseq2 library (Balioglu, 2023). We use a batch size of $64$ and a learning rate of 1e61e{-6} with a dropout rate of 0.0 and a $\beta$ value of 0.1 throughout the experiments. For each instruction, we generate 64 responses. These responses are then annotated with Athene-RM-8B to select preference pairs. Compared to human instructions, our synthetic instructions tend to be more complex, resulting in longer average response lengths, which can lead to length explosion. This occurs because, during DPO training, the evaluation judge often favors longer responses, potentially causing response lengths to increase over time (Yuan et al., 2024). To mitigate this issue, we adopted the approach outlined by Wu et al. (2024), which involves combining the reward score with length information to determine the preferred response. This method ensures that shorter responses are selected when scores are similar. We applied a length normalization coefficient of 0.2 for the length-normalized reward. This is applied for all methods, in each case constructing 5,000 DPO pairs.
@@
We also experiment with online DPO training by following the settings described by Lanchantin et al. (2025). We use the default sampling parameters (temperature=$1.0$, top-p=$1.0$) to generate exploration rollouts. We train models using the fairseq2 library (Balioglu, 2023), where model inference is performed with the vllm library (Kwon et al., 2023).
@@
To evaluate the helpfulness and quality of responses, we employ AlpacaEval 2.0 (Li et al., 2023b; Dubois et al., 2024) and Arena-Hard (Li et al., 2024b; a). These are robust instruction-following benchmarks that show a strong correlation with user preferences. Originally, AlpacaEval used GPT-4 Preview (11/06) as the judge, while Arena-Hard utilized GPT-4.1 for its leaderboard. However, since we do not have access to these specific OpenAI API versions, we conduct our tests using two alternative (and newer) judges: GPT-4-turbo and GPT-4o. For generating responses, we set the decoding temperature to 0.6 and the top-p to 0.9, aligning with the commonly used values of the seed model in our study. Our validation set, used for checkpoint selection, is based on a held-out set of 470 examples, comprising 253 validation examples from Li et al. (2023a) and 218 Evol-Test set examples from Xu et al. (2023).
@@ 5.1 Reasoning Tasks
In Table 1 where the Qwen3-4B-Base models are GRPO trained on Qwen3-4B generated instructions and targets, CoT-Self-Instruct achieves an average accuracy of 53.0%, outperforming Self-Instruct which yields 42.7% (both without data filtering). If we use Self-Instruct without CoT to generate instructions, but then use CoT to generate targets, we can achieve 49.5%, still inferior to CoT-Self-Instruct which uses CoT reasoning to generate both instructions and target answers. As shown in Appendix Table 8, similar trends are observed when training on Qwen3-4B-base, rather than Qwen3-4B, generated data. This highlights the importance of CoT thinking when generating reasoning instructions.
@@
Applying filtering methods to CoT-Self-Instruct and Self-Instruct improves both methods, despite the overall amount of training data decreasing, see Table 1. This suggests that it is better to have less high-quality synthetic data than more lower-quality data. However, we find that CoT-Self-Instruct still maintains its advantage over Self-Instruct, whichever filtering method is used. E.g. with Self-Consistency Filtering, Self-Instruct with CoT targets improves from 49.5% $\rightarrow$ 53.6%, while CoT-Self-Instruct improves from 53.0% $\rightarrow$ 55.1%. Similar findings are observed with RIP filtering as well. We find our proposed Answer-Consistency filtering, where answers are produced while generating an instruction with CoT, outperforms Self-Consistency filtering, which uses majority vote alone, with an average performance improvement from 55.1% $\rightarrow$ 57.2%.
Table 1: CoT-Self-Instruct results on reasoning tasks, compared to baselines, when fine-tuning Qwen3-4B-Base with GRPO. For Self-Instruct and CoT-Self-Instruct, the synthetic data (including targets) is constructed with Qwen3-4B. We report pass@1 averaged over 16 seeds. CoT-Self-Instruct generates synthetic data that outperforms existing training sets and the Self-Instruct method, particularly when applying our data filtering methods.
@@
s1k questions + (R1) gold label 893 68.6 18.5 51.3 40.1 44.6
OpenMathReasoning questions + gold label 10000 79.0 13.3 62.5 35.4 47.5
Self-Instruct questions + targets 5000 74.5 9.8 47.7 39.0 42.7
Self-Instruct questions + CoT-generated targets 5000 81.1 16.3 58.1 42.5 49.5
@@
+ Answer-Consistency Filter (more data) 10000 86.7 26.7 73.8 47.4 58.7
Table 2: CoT-Self-Instruct results on general instruction following tasks, comparing to baselines, fine-tuning LLama 3.1-8B-Instruct with offline and online DPO. CoT-Self-Instruct generates synthetic data that outperforms human-written instruction training sets and the Self-Instruct method, particularly when applying our data filtering methods. Both AlpacaEval 2 and ArenaHard are evaluated with two kinds of judge: GPT-4 Turbo and GPT-4o, with similar conclusions.
Training method AlpacaEval LC Winrate ArenaHard Score
GPT-4 Turbo GPT-4o GPT-4 Turbo GPT-4o Avg. $\uparrow$
@@
Human instructions (WildChat) DPO 49.1 43.0 52.7 42.6 46.8
@@
Human instructions (Wildchat) Online DPO 80.1 62.7 64.4 45.5 63.1
@@
High quality synthetic instructions generated by CoT-Self-Instruct significantly outperform seed instructions and other publicly available reasoning instructions
CoT-Self-Instruct outperforms s1k, as shown in Table 1, where the model trained on 2926 filtered examples using CoT-Self-Instruct achieve 57.2%. This is much higher than the 44.6% achieved with s1k instructions using R1 labels (and s1k results are even lower, 43.8%, with Qwen3-4B labels, see Appendix Table 5). Filtering CoT-Self-Instruct to the same training size as s1k yields 54.2%, still significantly higher, see Appendix Table 3. Our method also outperforms 10k OpenMath-Reasoning questions with gold labels, which gives 47.5%. Increasing the CoT-Self-Instruct with Answer-Consistency filtering data to 10k improves results further with an average of 58.7%. Overall, CoT-Self-Instruct with Answer-Consistency filter gives the best performance among all existing datasets or synthetic data construction methods tested.
Using other models or methods to generate targets yields similar conclusions
Our main experiments use Qwen3-4B-Base models GRPO trained on Qwen3-4B generated instructions and targets. We also experiment with several other settings. These include using Qwen3-4B-Base to generate targets (Appendix Table 8), Majority-vote to generate targets (Appendix Table 6), and Best-of-$K$ using a reward model to generate targets (Appendix Table 7). While each variant gives different overall maximum perfomance, in each case we see the same trend that CoT-Self-Instruct with Answer-Consistency is superior to Self-Instruct and other competing baselines.
5.2 General Instruction Following Tasks
@@
For non-reasoning tasks, allowing the model to create a plan beforehand with CoT-Self-Instruct also significantly enhances the quality of synthetic data, see Table 2. Averaged over AlpacaEval 2 and ArenaHard, CoT-Self-Instruct achieves an average of 53.9 vs. Self-Instruct’s 47.4, both without filtering and trained with DPO. We also observe that asking for longer CoT reasoning chains provides more gains than shorter CoTs (see Appendix Table 11), further emphasizing the need for reasoning even when producing synthetic data for non-verifiable general instruction following tasks.
@@
Applying the RIP filter to each method is proven to be effective across all types of synthetic generation methods tested. It boosts the CoT-Self-Instruct results from 53.9 $\rightarrow$ 54.7. RIP also improves Self-Instruct as well, from 47.4 $\rightarrow$ 49.1, but still underperforming CoT-Self-Instruct. We can also apply RIP filtering to human instructions from WildChat in a similar manner. In this case we actually see a larger boost, from 46.8 $\rightarrow$ 50.7. We attribute this to human data being relatively noisy compared to synthetic data, which can make filtering more important.
High quality synthetic instructions generated by CoT-Instruct significantly outperform human instructions
Our best performing DPO-trained model is achieved by using CoT-Self-Instruct with RIP data filtering, yielding 54.7. This outperforms Llama 3.1-8B-Instruct (27.1) or training on human instructions from WildChat with (46.8) or without RIP data filtering (50.7). We also performed experiments with online DPO, which improved results further. In that setting, human instructions from WildChat obtain 63.1 while CoT-Self-Instruct+RIP obtains 67.1. Overall, we find CoT-Self-Instruct with RIP filtering to yield the best performance over all existing datasets or synthetic data construction methods tested.
@@ 6 Conclusion
In this paper, we propose CoT-Self-Instruct, a synthetic data creation and curation pipeline that instructs LLMs to plan and reason to come up with new synthetic instructions given seed examples, and then filters them for quality, either using Answer-Consistency for verifiable tasks or RIP filtering when they are not verifiable. We show that applying our method improves models’ abilities in both reasoning and non-reasoning domains by creating high quality synthetic instructions for RL training, surpassing existing seed human-annotated instructions and public training sets on challenging benchmarks.